A bullet of mass 1g shot vertically up by a gun gives a force 2N for 0.001s.Calculate the maximum height attained by a bullet. Neglect air friction and take g=10ms-2. Solve this problem…………..

Grade:10

1 Answers

vikas askiitian expert
13 years ago

force = change in momentam/time

      2 =(mv-mu)/10-3 

           u=0 so

           v = 2m/s

now we have v2 = u2+2as

                  when bullet reaches maximum height then its velocity becomes 0..

                0 = u2-2gH

               H=u2/2g=0.2m